In a thermally insulated cylinder with an internal cross section S = 500 cm2 and a height l =




= 50 cm, there is a resistor with resistance R = 120 Ω. The cylinder is otherwise filled with air




at a temperature T0 = 20 ◦C and pressure p0 = 101 kPa, and the same kind of air surrounds




the cylinder. A current I = 200 mA flows through the resistor. A base of the cylinder breaks




away when pushed with a force exceeding F = 500 N. After what time does that happen?

1
Expert's answer
2021-11-21T17:25:06-0500

"F=(p-p_0)\cdot A" ,


"m=\rho V=1.29\cdot500\cdot10^{-4}\cdot0.5=0.03225\ (kg)"


"M=28.96\cdot10^{-3}\ kg/mole"


"\nu=m/M=0.03225/0.02896=1.114\ (mole)"




"p_0V_0/T_0=pV_0/T\to p=p_0T/T_0" or "F=(p_0T\/T_0-p_0)A"


"Q=\Delta U\to Q=\frac{i}{2}\nu R(T-T_0)\to I^2rt=\frac{i}{2}\nu R(T-T_0)"


"T=\frac{2I^2rt}{i\nu R}+T_0"


"F=(p_0(\frac{2I^2rt}{i\nu R}+T_0)/T_0-p_0)A\to t=\frac{T_0i\nu RF}{2I^2rp_0A}="


"=\frac{293\cdot5\cdot 1.114\cdot8.31\cdot500}{2\cdot 0.2^2\cdot 120\cdot101000\cdot 0.05}=140\ (s)" . Answer